Hayhayz (hayhayz):

In your lab, a substance's temperature has been observed to follow the function T(x) = (x − 4)3 + 6. The turning point of the graph is where the substance changes from a liquid to a gas. Using complete sentences in your written answer, explain to your fellow scientists how to find the turning point of this function. Hint: The turning point of the graph is similar to the vertex of a quadratic function. @dan815

OpenStudy (dan815):

\[T(x) = (x − 4)^3 + 6\]

Hayhayz (hayhayz):

Yes that's the equation

OpenStudy (dan815):

okay so at a turning point the slope =0

OpenStudy (dan815):

Can you tell me what the derivative of this function is

Hayhayz (hayhayz):

Uhh i do not know

OpenStudy (dan815):

have you learnt how to take derivatives yet

OpenStudy (dan815):

if u havent then you have to know what y=x^3 graph looks like

Hayhayz (hayhayz):

No.. What's a derivative

OpenStudy (dan815):

oh okay, then its fine, dont worry about it

OpenStudy (dan815):

What does y=x^3 look like

Hayhayz (hayhayz):

 something like that?

OpenStudy (dan815):

okay yep, so where is the turning point for this graph

Hayhayz (hayhayz):

X=0

OpenStudy (dan815):

okay so now when we apply transformations

OpenStudy (dan815):

this point will be shifted

OpenStudy (dan815):

right now the turning point is at (0,0)

OpenStudy (dan815):

see each transformation one at a time

OpenStudy (dan815):

y=x^3+6 <---- if we apply only 6, so everywhere you see x^3 we add 6 to all of it

OpenStudy (dan815):